OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 56°, a = 12, b = 14

OpenStudy (anonymous):

I'm having trouble solving for c, I've already been able to solve for both B1 and B2 as well as C2 and C1.

OpenStudy (anonymous):

Eyeball this http://hotmath.com/hotmath_help/topics/law-of-sines.html and take a stab at it.

OpenStudy (anonymous):

I'm assuming side \(a\) is the side opposite angle \(A\), and likewise for \(b/B\) and \(c/C\). The law of sines gives you \[\begin{align*}\frac{\sin A}{a}&=\frac{\sin B}{b}\\ \frac{\sin56^\circ}{12}&=\frac{\sin B}{14} \end{align*}\] Solve for angle \(B\). The interior angles of a triangle add up to 180°, which means \[A+B+C=180^\circ\] You have \(A\) already, you'll have \(B\) from the law of sines equation, so solve for \(C\). Another application of law of sines to find the remaining side \(c\): \[\begin{align*}\frac{\sin A}{a}&=\frac{\sin C}{c}\\ \frac{\sin56^\circ}{12}&=\frac{\sin C}{c} \end{align*}\]

OpenStudy (anonymous):

I've been using that same formula as said in your link, but get a completely different answer. \[12\div \sin56 = c \div \sin48.7\]

OpenStudy (agent0smith):

Show your work, what you got?

OpenStudy (anonymous):

A = 56 B1 = 75.3 C1 = 48.7 B2 = 104.7 C2 = 19.3 a = 12 b = 14 @agent0smith

OpenStudy (anonymous):

The answers are: B = 75.3°, C = 48.7°, c = 10.9; B = 104.7°, C = 19.3°, c = 4.8 B = 14.7°, C = 109.3°, c = 13.7; B = 165.3°, C = 70.7°, c = 13.7 B = 75.3°, C = 48.7°, c = 13.2; B = 104.7°, C = 19.3°, c = 13.2 B = 14.7°, C = 109.3°, c = 10.5; B = 165.3°, C = 70.7°, c = 10.5